- Does a reaction occur when aqueous solutions of magnesium sulfate and copper(II) nitrate are combined?

- Does a reaction occur when aqueous solutions of barium hydroxide and cobalt(II) nitrate are combined?

- Write the net ionic equation for the following molecular equation.

MgCl2(aq) + (NH4)2CO3(aq) = MgCO3(s) + 2NH4Cl(aq)

(Use the lowest possible coefficients. Use the pull-down boxes to specify states such as (aq) or (s). If a box is not needed, leave it blank.)

The following molecular equation represents the reaction that occurs when aqueous solutions of lead(II) nitrate and barium bromide are combined.

Write the balanced net ionic equation for the reaction.

Pb(NO3)2(aq) + BaBr2(aq) = PbBr2(s) + Ba(NO3)2(aq)

(Use the lowest possible coefficients. Use the pull-down boxes to specify states such as (aq) or (s). If a box is not needed, leave it blank.)
Answer #1

1. There is no reaction between MgSO4 and Cu(NO3)2.

2. We have the following reaction:

Ba(OH)2(aq)+ Co(NO3)2(aq) Ba(NO3)2(aq) + Co(OH)2(s)

3. The given equation is balanced and correct.

MgCl2(aq) + (NH4)2CO3(aq) = MgCO3(s) + 2 NH4Cl(aq)

4. Balanced equation:

Pb(NO3)2(aq) + BaBr2(aq) = PbBr2(s) + Ba(NO3)2(aq)